Shift Summary for owl shift, 11/04/99

John Shields (jshields@ksera)
Thu, 4 Nov 1999 07:07:05 -0600

Shift Summary for owl shift, 11/04/99

On Shift: J. Shields, J. Ladue

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

tevatron still down (cyro problems).
========================================================================
Runs ended this shift:
Run 14902 - Cosmics

Tapes brought to FCC this shift:
(including tapes for online split)

none
========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:

basically very, very, very quiet. The silence was broken only by a
few scattered ktevana errors, a csi gain drift complaint, and a
pipeline error.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

02:11 MCR calls & says there are some cyro problems.
As a result, the chances of getting beam before the
scheduled 7 am shutdown are extremely slim.

05:20 triggers hang and refuse to be pacified by kfrends
and a few other common little quick-fixes. Shift
crew eventually cops out and simply starts a new
cosmic run. Problem solved.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift